### v1

# Towards Explainable Bilingual Multimodal Misinformation Detection and Localization

Abstract
The increasing realism of multimodal content has made misinformation more subtle and more complex to detect, especially in news media, where images are frequently paired with bilingual subtitles (e.g., Chinese-English). Such content often includes localized image edits and cross-lingual inconsistencies that jointly distort meaning while remaining superficially plausible. We introduce BiMi, a bilingual multimodal framework that jointly performs region-level localization, cross-modal and cross-lingual consistency detection, and natural language explanation for misinformation analysis. To support generalization, BiMi integrates an online retrieval module that supplements model reasoning with up-to-date external context. We further release BiMiBench, a large-scale and comprehensive benchmark constructed by systematically editing real news images and subtitles, comprising 104,000 samples with realistic manipulations across visual and linguistic modalities. To enhance interpretability, we apply Group Relative Policy Optimization (GRPO) to improve explanation quality, marking the first use of GRPO in this domain. Extensive experiments demonstrate that BiMi outperforms strong baselines by up to +8.9 in classification accuracy, +15.9 in localization accuracy, and +2.5 in explanation BERTScore, advancing state-of-the-art performance in realistic, multilingual misinformation detection. Code, models, and datasets are avaliable at https://github.com/RaffeLegend/misinformation.
1 Introduction
The rapid advancement of generative models [1, 2, 3, 4, 5] has drastically lowered the barrier to producing highly realistic multimodal content [6]. While enabling creative applications, this progress has also escalated the proliferation of misinformation that intricately blends visual and textual elements [7, 8]. A particularly insidious form emerges in news media, where images are frequently overlaid with bilingual subtitles (e.g., Chinese-English). Here, subtle visual tampering and cross-lingual inconsistencies are exploited to distort facts while maintaining surface-level plausibility. This becomes especially dangerous for monolingual audiences, who may only comprehend one language and thus remain oblivious to manipulations in the other. A stark example occurred during early 2020, when a CDC report confirming the first U.S. case of COVID-19 community transmission [^1] https://edition.cnn.com/2020/02/27/health/us-cases-coronavirus-community-transmission/index.html was mistranslated by some Chinese social media as asserting that the virus originated in the United States. This misleading bilingual manipulation fueled significant public misunderstanding and geopolitical tensions. Such cases illustrate how multimodal misinformation increasingly leverages localized image edits and asymmetric subtitle translations to manipulate public perception across language communities. Addressing such sophisticated attacks demands more than conventional binary misinformation detection. It requires systems capable of region-level localization of manipulations, identification of cross-modal and cross-lingual inconsistencies, and generation of coherent, cross-lingual natural language explanations—capabilities that remain underexplored.
Despite encouraging progress [7, 8, 9, 10, 11], existing systems exhibit three critical limitations. First, most rely on global image-text matching, lacking the fine-grained region-level localization necessary to expose subtle manipulations. Second, their explanations are often superficial and generic, providing little actionable insight into why content is misleading. Third, they neglect the modeling of bilingual subtitle inconsistencies, overlooking the semantic divergence arising from intentionally misleading translations. These shortcomings underline the need for a unified framework that reasons at a fine granularity across visual and bilingual textual modalities.
Figure 1: Comparison of traditional vs. bilingual misinformation detection tasks. Traditional tasks focus on visual-text consistency with limited outputs (left). Our setting uses tampered news images and bilingual subtitles, enabling richer outputs including region localization, cross-modal consistency, and explanation (right). Red indicates error, green indicates correctness. Best viewed in color.
However, building such a framework is technically challenging. Region localization demands precise visual-textual alignment that is beyond the capacity of most pretrained MLLMs. Generating faithful explanations in bilingual, multimodal settings requires accurate identification and articulation of inconsistencies across modalities and languages. Moreover, cross-lingual reasoning introduces unique complexities, as even fluent translations may introduce subtle but impactful semantic shifts that obscure the misinformation signal. These challenges are further exacerbated in the news domain, characterized by rapidly evolving, diverse, and context-dependent content.
To tackle these challenges, we introduce BiMi, the first bilingual multimodal misinformation detection framework that seamlessly integrates region-level manipulation localization, cross-modal and cross-lingual consistency detection, and natural language explanation generation. An overview of our task and outputs is illustrated in Figure 1. To enhance generalization to emerging events, BiMi is equipped with an online retrieval module that augments model reasoning with real-time external knowledge. Furthermore, to facilitate systematic evaluation, we construct BiMiBench, the first large-scale benchmark for this setting, comprising 104,000 samples with realistic manipulations of news images and bilingual subtitles. Through extensive experiments, BiMi achieves state-of-the-art performance, surpassing strong baselines by +8.9 in classification accuracy, +15.9 in localization accuracy, and +2.5 in explanation BERTScore.
Our contributions are summarized as follows:
  • We release BiMiBench, the first large-scale benchmark for bilingual multimodal misinformation detection, containing 104K samples with fine-grained image manipulations and bilingual subtitle inconsistencies.
  • We propose BiMi, a unified framework that detects image-subtitle misinformation via region localization, cross-modal consistency detection, and explanation generation, addressing the limitations of prior works.
  • We introduce the first application of Group Relative Policy Optimization (GRPO) [12] in this domain, enhancing explanation quality in multimodal, bilingual scenarios.
  • Our framework achieves new state-of-the-art performance on BiMiBench, substantially improving classification, localization, and explanation metrics over competitive baselines.
2 Related Work
2.1 Misinformation Detection
Datasets. Early misinformation detection datasets can be grouped into three main categories. Text-only datasets: FEVER [13] and FakeNewsNet [14] focus exclusively on textual fact verification and lack any visual modality, limiting their applicability to multimodal scenarios. Basic multimodal datasets: Fakeddit [15], MAIM [16], and EMU [17] introduce image-text pairs, but are typically limited to monolingual (English) content, coarse visual edits, and offer no explicit supervision for localization or explanation. Semantically aligned multimodal datasets: More recent benchmarks such as NewsCLIPpings [18], COSMOS [19], MMFakeBench [8] and DGM4 [9] move toward finer-grained multimodal reasoning. However, they remain focused on caption-style inputs and overlook subtitle-level or cross-lingual inconsistencies.
To address these limitations, we introduce BiMiBench, a large-scale benchmark curated from real-world news content that captures both visual and textual manipulations across five fine-grained misinformation types. As shown in Table 2.1, BiMiBench is the only benchmark that simultaneously supports bilingual subtitles, region-level localization, and both visual and textual tampering, offering the richest annotation scope among existing datasets. To evaluate the quality of visual manipulations, Figure 2 compares BiMiBench with recent benchmarks using a perceptual realism analysis across seven dimensions (e.g., content, composition, color), demonstrating that BiMiBench achieves significantly higher realism and diversity.
Table 1: Comparison of misinformation datasets. BiMiBench uniquely supports manipulations based on bilingual subtitles and localized content.
Dataset Multi- Sub- Bi- Textual Visual Local- News Mask
modal title lingual Manip. Manip. ization Domain
FEVER [13]
FakeNewsNet [14]
Fakeddit [15]
MAIM [16]
EMU [17]
NewsCLIPpings [18]
COSMOS [19]
DGM4 [9]
MMFakeBench [8]
BiMiBench (Ours)
Figure 2: Quality comparison of manipulated images across datasets, by the method of [20].
Methods. Current methods for multimodal misinformation detection typically focus on aligning image and text pairs, with an emphasis on out-of-context detection [21, 22, 9, 23, 24, 25]. NewsCLIPpings [18] adopts CLIP-based matching for image-caption congruence. COSMOS [19] employs self-supervised contrastive learning on out-of-context pairs but lacks supervision for fine-grained or localized manipulations. SNIFFER [7] and EMU [17] improve explainability via MLLMs and intent reasoning, yet rely on clean, monolingual captions. HAMMER [9] localizes manipulations through grounding or matching, but does not address subtitle-level or multilingual inconsistencies. MMD-Agent [8] broadens evaluation with mixed-source samples, though without explicit modeling of bilingual or region-specific manipulations. CroMe [26] introduces cross-modal metric learning over image, text, and fused representations, but remains limited to English captions. In contrast, we are the first to detect misinformation involving tampered images and bilingual subtitles, with support for localization, cross-lingual reasoning, and explanation.
2.2 Multi-modal Large Language Models
Model. Multimodal large language models (MLLMs) [27, 4, 28, 29, 30] integrate vision encoders with pretrained language models to enable unified visual-text reasoning. Flamingo [31] introduces prompt-based multimodal learning, showing strong few-shot results on tasks like VQA. BLIP-2 [28], MiniGPT-4 [29], and LLaVA series [30, 32, 33] further enhance visual instruction tuning, aligning modalities through instruction-following data to support tasks such as VQA, captioning, and dialogue. Recent models like InternVL3 [27], Qwen3-VL [34], and Gemini [35] improve generalization, multilinguality, and open-ended tool use. Despite these advances, most MLLMs focus on grounded understanding or generation and lack supervision for detecting fine-grained cross-modal inconsistencies—critical for misinformation detection [36, 37]. To improve efficiency, many lightweight MLLMs have emerged, such as Gemma [38]; in our work, we adopt Gemma 3 [39] as the base model to balance performance and scalability.
GRPO. Reinforcement learning (RL) has been shown to enhance reasoning in LLMs [40], with DeepSeek-R1 [5, 12] introducing Group Relative Policy Optimization (GRPO) to further improve explanation-driven multimodal performance. GRPO optimizes the relative ranking of outputs within grouped samples, enabling fine-grained discrimination between aligned and misaligned multimodal inputs. However, its potential remains underexplored in misinformation detection tasks involving subtle visual edits and bilingual subtitle inconsistencies.
Prior work [7, 8, 9] rarely tackles misinformation that involves both manipulated images and bilingual subtitles, and often lacks the ability to localize edits, reason across modalities, and adapt to new events. To address these gaps, we propose a unified framework that integrates bilingual visual-text reasoning, manipulation localization, explanation generation, and retrieval-based adaptation.
3 BiMiBench: A Benchmark for Bilingual Multimodal Misinformation
Existing benchmarks for multimodal misinformation primarily focus on monolingual captions or synthetic mismatches, lacking realism, cross-lingual scope, and fine-grained supervision. In real-world scenarios, misinformation often involves tampered images with bilingual (Chinese-English) subtitles, where inconsistencies may occur in any modality. To address this gap, we introduce BiMiBench, the first benchmark that supports not only bilingual multimodal reasoning but also manipulation, localization, and natural language explanation, enabling comprehensive evaluation in realistic settings.
3.1 Dataset Construction
Details. BiMiBench contains 104,000 samples derived from real-world news content. Specifically, we construct our dataset based on the VisualNews [41] dataset and select the portion sourced from the Washington Post to ensure factual grounding and linguistic consistency. Each sample includes an image paired with Chinese-English subtitles, with misinformation introduced through localized image editing, Chinese subtitle modification, or English subtitle modification. Approximately 80% of the samples contain manipulated content in at least one modality, while the remaining 20% are clean control examples. Image resolutions range from 640×480 to 1024×768 pixels, and subtitles are provided in both English and Chinese to support bilingual misinformation detection. BiMiBench captures cross-modal inconsistencies, offering a challenging testbed for evaluating multimodal and bilingual reasoning capabilities.
Construction. Our benchmark is constructed based on VisualNews, a large-scale image-text news dataset.
Figure 3: The data generation workflow used in constructing the BiMiBench benchmark.
To generate realistic misinformation samples, we design a multi-step data generation pipeline. Let $\mathcal{D}_{\text{orig}}=\{(I_{i},C_{i})\}_{i=1}^{N}$ denote the original dataset from VisualNews, where $I_{i}$ is an image and $C_{i}$ is its associated English caption.
Step 1: Textual Manipulation. Given an image-caption pair $(I_{i},C_{i})$ from VisualNews [41], we use the Gemma 3 [39] model $f_{\text{Gemma}}$ to generate a manipulated caption $\tilde{C}_{i}=f_{\text{Gemma}}(I_{i},C_{i},\text{PROMPT})$, introducing factual inconsistencies while preserving contextual plausibility.
Step 2. Bilingual Translation. We translate the original and manipulated English captions into Chinese using a translation API [42] $\mathcal{T}$, yielding bilingual subtitle pairs: $C^{\text{zh}}_{i}=\mathcal{T}(C_{i}),\tilde{C}^{\text{zh}}_{i}=\mathcal{T}(\tilde{C}_{i})$.
Step 3. Semantic Segmentation. To enable targeted visual editing, we use an instruction-based segmentation model LISA [43]:$f_{\text{LISA}}$ to extract object masks based on the manipulated caption: $M_{i}=f_{\text{LISA}}(I_{i},\tilde{C}_{i},\text{PROMPT})$, where $M_{i}$ highlights regions corresponding to entities in $\tilde{C}_{i}$.
Step 4. Visual Manipulation. We apply a visual manipulation function $\mathcal{V}$ to perform localized editing on the original image using the object masks: $\tilde{I}_{i}=\mathcal{V}(I_{i},M_{i})$, where the edited image $\tilde{I}_{i}$ semantically aligns with the manipulated caption $\tilde{C}_{i}$. To ensure visual diversity and realism, we adopt a mix of recent state-of-the-art image editing techniques, including FLUX [44], VAR [45], and SDXL [46].
Step 5. Label Assignment. For each final sample, we randomly select real or manipulated versions of the image and subtitles to construct a multimodal example: $S_{i}^{*}=S\left(I_{i}^{*},C_{i}^{\text{en}},L_{i}^{\text{zh}}\right),I_{i}^{*}\in\{I_{i},\tilde{I}_{i}\},C_{i}^{\text{en}}\in\{C_{i},\tilde{C}_{i}\},L_{i}^{\text{zh}}\in\{C^{\text{zh}}_{i},\tilde{C}^{\text{zh}}_{i}\}$, where misinformation may appear in any modality individually or in combination.
Table 2: BiMiBench Category Definitions.
Category Definition
All Consistent $I_{i}^{*}=I_{i},C_{i}^{\text{en}}=C_{i},L_{i}^{\text{zh}}=C_{i}^{\text{zh}}$
Image Manipulated $I_{i}^{*}=\tilde{I}_{i},C_{i}^{\text{en}}\neq\tilde{C}_{i}$ or $L_{i}^{\text{zh}}\neq\tilde{C}_{i}^{\text{zh}}$
Both Misaligned $I_{i}^{*}=I_{i},C_{i}^{\text{en}}=\tilde{C}_{i},L_{i}^{\text{zh}}=\tilde{C}_{i}^{\text{zh}}$
Chinese Misaligned $I_{i}^{*}=I_{i},C_{i}^{\text{en}}=C_{i},L_{i}^{\text{zh}}=\tilde{C}_{i}^{\text{zh}}$
English Misaligned $I_{i}^{*}=I_{i},C_{i}^{\text{en}}=\tilde{C}_{i},L_{i}^{\text{zh}}=C_{i}^{\text{zh}}$
All Inconsistent $I_{i}^{*}=\tilde{I}_{i},C_{i}^{\text{en}}=\tilde{C}_{i},L_{i}^{\text{zh}}=\tilde{C}_{i}^{\text{zh}}$
Each sample in BiMiBench is assigned to one of six mutually exclusive categories based on image manipulation and semantic alignment between the image $I_{i}^{*}$, English subtitle $C_{i}^{\text{en}}$, and Chinese subtitle $L_{i}^{\text{zh}}$. The categories cover typical multimodal misinformation patterns: (1) All Consistent: Image and both subtitles are consistent. (2) Image Manipulated: The image is tampered, and at least one subtitle does not align with it. (3) Both Misaligned: Image is real, but both subtitles are misleading. (4) Chinese Misaligned: Image is real, English subtitle is faithful, Chinese is misleading. (5) English Misaligned: Image is real, Chinese subtitle is faithful, English is misleading. (6) All Inconsistent: Image is tampered; both subtitles are misleading. Table 2 provides formal definitions for each class. All samples are manually reviewed for annotation quality and consistency. This design enables structured supervision of complex multimodal misinformation scenarios. Additional examples and detailed analysis of BiMiBench are provided in the Appendix.
4 BiMi: A Bilingual Multimodal Misinformation Detection Framework
Real-world multimodal misinformation often involves image edits and inconsistencies between bilingual subtitles. Existing models struggle to detect such cross-modal manipulations and lack adaptability to emerging events. We propose BiMi, a bilingual multimodal framework that localizes manipulated regions, detects cross-modal inconsistencies, and generates explanations. To enhance generalization, BiMi integrates an online retrieval module that provides real-time external context.
Figure 4: The overview of the training strategy. Three stages: domain alignment on news data, instruction tuning with task-specific prompts, and GRPO optimization with structured rewards. Each stage builds upon the previous one, progressively enhancing task performance.
4.1 Framework
Modeling. We adopt Gemma 3 [39] as the pre-trained model of BiMi due to its strong multilingual understanding and cross-modal reasoning capabilities. Gemma 3 is a multilingual MLLM with strong vision-language alignment and cross-lingual reasoning, making it well-suited for detecting subtle manipulations across images and bilingual subtitles. In our framework, the input image is overlaid with Chinese and English subtitles to form a unified visual input, which is then processed by a vision encoder to extract patch-level embeddings. These representations are fused through learned attention layers to form a unified multimodal context. This design allows the model to jointly reason over visual and bilingual textual cues, making it well-suited for detecting subtle manipulations such as object removal, semantic shifts, and cross-lingual inconsistencies.
Retrieval Module. To improve adaptability to emerging misinformation, BiMi integrates a retrieval module at inference time. We first extract Chinese and English subtitles from the image using OCR, then query the Google Search API [42] and collect the top-3 results. Retrieved texts are appended to the input prompt, providing up-to-date context that helps the model assess semantic consistency. This retrieval-augmented setup enables BiMi to generalize better to previously unseen or rapidly evolving misinformation cases.
4.2 Multi-stage Training
To effectively adapt the pretrained Gemma to the task of multilingual multimodal misinformation detection, we adopt a three-stage training strategy aimed at progressively aligning the model with the news domain, task-specific instructions, and high-quality reasoning objectives. Figure 4 shows the overview of the training strategy.
Stage 1: Domain Alignment. To adapt the model to the linguistic and visual characteristics of the news domain, we perform instruction-based tuning using image-caption pairs from VisualNews [41]. Each instance consists of an instruction prompt (e.g., Describe this image in English/Chinese) and a corresponding caption in either English or Chinese. We use captions from the Washington Post subset and translate each English caption into Chinese to ensure cross-lingual coverage. The model learns to generate captions conditioned on both the image and the prompt, thereby establishing strong visual-textual grounding across both languages. This stage enhances the model’s ability to handle real-world journalistic content and supports bilingual reasoning in downstream misinformation detection tasks.
Stage 2: Instruction Tuning. We fine-tune the model with single-turn instruction-following data tailored for multimodal misinformation detection. Each sample includes an image $X_{v}$, a bilingual subtitle pair, and a prompt $X_{\text{q}}$ that asks the model to perform multiple sub-tasks: detect manipulation, assess cross-modal consistency, and explain the decision. The assistant response $X_{\text{a}}$ includes structured answers and a natural language explanation. We follow a unified sequence format:
[…]
where $X_{\text{q}}$ contains the natural language instruction and bilingual subtitles, and $X_{\text{a}}$ consists of three binary labels and a free-form explanation, formatted as:
[…]
The model is trained to generate $X_{\text{a}}$ conditioned on $X_{\text{q}}$ and $X_{v}$, by maximizing the likelihood:
$P(X_{\text{a}}\mid X_{\text{q}},X_{v})=\prod_{i=1}^{L}P_{\theta}(x_{i}\mid x_{<i})$ (1)
where $x_{i}$ denotes the $i$-th token in the assistant response. Only the tokens in $X_{\text{a}}$ are used to compute the loss. This tuning step encourages the model to align with the structure and reasoning required for fine-grained misinformation detection.
Stage 3: GRPO-based Reasoning Optimization. To improve the model’s reasoning quality and explanation consistency, we apply Group Relative Policy Optimization (GRPO) [12], a reinforcement learning strategy that adjusts the policy by comparing outputs within a batch rather than relying on absolute rewards alone.
Given a question $q$, GRPO samples $N$ candidate responses $\{r_{1},r_{2},\ldots,r_{N}\}$ from the policy $\pi_{\theta}$ and evaluates each response $o_{i}$ using a reward function $R\left(q,o_{i}\right)$, which measures the quality of the candidate in the context of the given question. GRPO encourages the model to generate responses with higher advantages within the group by updating the policy $\pi_{\theta}$ using the following objective:
$\mathcal{J}_{\mathrm{GRPO}}(\theta)=\mathbb{E}\left[{\{o_{i}\}_{i=1}^{N}\sim\pi_{\theta_{\text{old}}}(q)}\right]\frac{1}{N}\sum_{i=1}^{N}\left\{\min\left[s_{1}\cdot A_{i},\;s_{2}\cdot A_{i}\right]-\beta\mathbb{D}_{\mathrm{KL}}\left(\pi_{\theta}\|\pi_{\mathrm{ref}}\right)\right\}$ (2)
$A_{i}$ $=\frac{r_{i}-\mathrm{mean}\{r_{1},r_{2},\ldots,r_{N}\}}{\mathrm{std}\{r_{1},r_{2},\ldots,r_{N}\}},\quad s_{1}=\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)},\quad s_{2}=\mathrm{clip}\left(s_{1},\;1+\epsilon,\;1-\epsilon\right),$ (3)
where $A_{i}$ represents the advantage of the candidate response $o_{i}$ relative to other sampled responses. Following DeepSeek-R1, we use both format and accuracy reward.
Reward function. To optimize the reasoning ability and explanation quality of BiMi during the final training stage, we design a composite reward function tailored for GRPO. In our setting, we focus on three core tasks: misinformation classification, tampered region localization, and natural language explanation generation.
Format reward. To enforce structured outputs, we define a format reward $R_{\text{format}}$ that checks whether the model response follows a predefined format, using <answer></answer> tags for predictions and an optional <think></think> section for explanations. This ensures consistency across outputs and simplifies automatic evaluation. We use regular expressions for validation: $R_{\text{format}}=\begin{cases}0,&\text{if output matches expected format}\\ 1,&\text{otherwise}\end{cases}$.
Localization reward. We use an IoU-based reward $R_{\text{Loc}}$ to assess the alignment between predicted and ground-truth tampered regions. To simplify computation, both are represented as bounding boxes. The reward is defined as: $R_{\text{Loc}}=\frac{|\mathcal{M}_{\text{pred}}\cap\mathcal{M}_{\text{gt}}|}{|\mathcal{M}_{\text{pred}}\cup\mathcal{M}_{\text{gt}}|}$, where $\mathcal{M}_{\text{pred}}$ and $\mathcal{M}_{\text{gt}}$ denote the predicted and ground-truth bounding boxes, respectively.
Classification reward. To encourage accurate decision-making in multimodal misinformation detection, we define a classification reward $R_{\text{cls}}$ that compares the model’s predicted label with the ground-truth annotation. This applies to tasks such as detecting whether the image is manipulated or whether the subtitles are semantically consistent. $R_{\text{cls}}=\begin{cases}1,&\text{if }C_{\text{pred}}=C_{\text{gt}}\\ 0,&\text{otherwise}\end{cases}$.
The final reward combines all task-specific objectives, including formatting, classification, localization: $R_{\text{total}}=R_{\text{format}}+R_{\text{cls}}+R_{\text{loc}}$. This unified reward encourages the model to generate structured, accurate, and interpretable predictions across modalities.
This progressive training strategy equips BiMi with the capability to perform fine-grained classification, localization, and explanation, and is designed to support generalization to challenging, real-world misinformation cases.
5 Experiments
5.1 Experimental Setup
We evaluate BiMi on BiMiBench and MMFakeBench [8]. On BiMiBench, the model performs six-class classification, tampered region localization, and explanation generation. For MMFakeBench, which lacks subtitle information, we follow its four-class setting and provide text context via prompts.
Implementation Details. We use the Gemma 3 as the LLM, applying our three-stage post-training pipeline. All experiments are run on 48GB V100 GPUs with a batch size of 16, learning rate $1\times 10^{-5}$, and frozen vision encoder. GRPO is applied for two additional epochs. We design structured instruction prompts to guide the model in tasks, as detailed in Table 3.
Please analyze the given image containing both Chinese and English subtitles and complete the following three tasks: (1) Classification Task: classify the alignment between the image and the subtitles into one of the following six categories: "all consistent", "image manipulated", "both subtitles misaligned with image", "only English aligned", "only Chinese aligned", "all inconsistent". (2) Manipulation Detection: if the image has been manipulated, return one or more bounding boxes for the manipulated regions in the format: {"bbox":[x_min, y_min, x_max, y_max]}. If no manipulation is found, return an empty list. (3) Decision Explanation: briefly explain your thinking before the classification and any detected regions. Return your output using the following format, wrapped in tags:
<think>Your explanation here.</think>
<answer>{"classification": "result", "region": [{"bbox": [x_min, y_min, x_max, y_max]}]}</answer>
Table 3: Prompt templates used in BiMi for multimodal misinformation detection
Baselines. We compare BiMi against a range of strong multimodal baselines, including InternVL3 [27], Qwen3 [47], LLaVA-1.6 [33], and LLama3.1 [48], which represent leading approaches in vision-language modeling. We also include specialized misinformation detection systems such as SNIFFER [7] and MMD-Agent [8], designed for out-of-context and multimodal misinformation. All models are evaluated using the same inputs: the image with overlaid Chinese-English subtitles and a task-specific prompt. For ablation, we test BiMi variants without GRPO, without retrieval, and with alternative MLLM backbones.
Evaluation Metrics. We report Accuracy (ACC) and F1 score for classification, IoU for localization. Inspired by [30], we evaluate explanation quality via BERTScore [49] between response and reference explanations generated by prompting GPT-4o [50] with labels and manually verified for correctness.
Table 4: Main results on BiMiBench. BiMi achieves the best performance across all metrics. Bold and underline denote best and second-best results.
Method LLM Res. ACC F1 IoU BERT
InternVL3 [27] Qwen2.5-7B 224 20.85 18.42 7.23 71.82
InternVL3 [27] Qwen2.5-14B 224 28.47 26.71 14.38 76.54
Qwen3 [47] Qwen3-8B 224 21.43 14.96 6.28 73.92
LLaVA-1.6 [33] Vicuna-7B 224 22.91 20.74 8.91 75.77
LLama3.1 [48] Llama-3.1-8B 224 24.61 21.94 5.27 72.91
Gemma3 [48] Gemma3-4B-IT 336 17.48 11.73 4.29 70.84
SNIFFER [7] Vicuna-13B 224 37.90 33.63 80.41
MMD-Agent [8] Qwen2.5-7B 224 34.78 29.48
BiMi Gemma3-4B-IT 336 46.80 42.79 30.24 82.90
5.2 Main Results
Table 5: MMFB results.
Method ACC F1
InternVL3 42.69 34.80
Qwen3 44.82 39.49
LLaVA-1.6 30.09 24.31
LLama3.1 32.67 28.30
Gemma3 27.19 22.04
SNIFFER 67.49 61.44
MMD-Agent 62.78 52.83
BiMi 70.94 62.49
Performance Comparison. We compare BiMi with strong baselines across three subtasks on the BiMiBench test set: (1) misinformation detection, (2) tampered region localization, and (3) explanation generation. Results are summarized in Table 4. BiMi significantly outperforms all baselines in both all tasks. In particular, it achieves a +8.9 accuracy gain in image-subtitle consistency, a +15.9 localization accuracy, and a +2.5 BERTScore improvement in explanation generation over the strongest baseline. BiMi’s performance also benefits from the GRPO reward design, which improves explanation specificity and localization precision. The results on MMFakeBench (4-class classification task) are shown in Table 5.
Ablation Study. We investigate the contributions of each stage in our pipeline using an ablation study. (i) Domain Alignment. Removing this stage reduces accuracy by 2.35 and BERTScore by 4.48, indicating its role in grounding visual-textual reasoning on news content. (ii) Instruction Tuning. Eliminating instruction tuning leads to a drastic performance drop, accuracy falls by 31.18 and BERTScore by 12.06, highlighting its importance in enabling structured bilingual understanding. (iii) GRPO. Without GRPO, the model fails to generate specific explanations tied to image or
Table 6: Table: Ablation results (%). Red text shows a performance drop from the best variant.
Design Accuracy BERTScore
Full model 48.60 82.90
w/o Domain Align 46.25 (-2.35) 78.42 (-4.48)
w/o Instr. Tune 17.42 (-31.18) 70.84 (-12.06)
w/o GRPO 42.90 (-5.70) 74.81 (-8.09)
Answer First 46.66 (-1.94) 81.79 (-1.11)
subtitle content, reducing accuracy by 5.7 and BERTScore by 8.09. (iv) Output Order. Reversing the output order (classify before explain) slightly lowers accuracy (–1.94) and BERTScore (–1.11), suggesting that generating explanations first facilitates more coherent reasoning. Table 6 summarizes the results. These findings confirm that all components are essential, with instruction tuning and GRPO being particularly critical for boosting accuracy and explanation quality.
5.3 Explainability Analysis
Beyond accurate classification and localization, generating high-quality natural language explanations is crucial for interpretable misinformation detection. We analyze explanation ability through (1) qualitative comparisons, (2) the influence of the retrieval module, and (3) evaluation under bilingual.
Explanation Quality. To understand BiMi’s reasoning ability, we compare its output with InternVL3 [27] on a representative case (Figure 5). Both models recognize that the subtitles are semantically aligned, showing bilingual understanding. However, InternVL3 relies on external factual cues and overlooks the mismatch between image and text. In contrast, BiMi grounds its decision in the image, correctly identifies the manipulation, and localizes the tampered region. It also provides a clearer, step-by-step explanation, showing stronger visual grounding and more faithful reasoning.
Figure 5: Comparison of explanations from InternVL3 (middle) and our model (right). The top left is input. The bottom left is the original sample. Note: Due to space constraints, some model responses are truncated. Full examples and additional results are provided in the Appendix.
Effect of Retrieval. We analyze a real-world social media post with subtle bilingual inconsistency (Figure 6). The removed "student" in the Chinese translation alters the meaning. Without retrieval, the model misclassifies due to a lack of context. The retrieval result provides a key clue for classification. BiMi correctly identifies the inconsistency, showing its benefit for nuanced cross-lingual reasoning.
Table 7: Accuracy and BERTScore for each language variant (4-category classification).
Input Variant ACC BERTScore
CN-only 72.32
EN-only 82.48 83.44
Bilingual multimodal capability. To evaluate bilingual understanding, we construct a variant of the test set where the model receives only one subtitle, Chinese or English. This reduces the task to a 4-category classification. By ignoring cross-lingual consistency, the single-language setting reduces reasoning difficulty. As shown in Table 7, the model performs better with English-only inputs, likely due to the predominance of English data during training. This highlights a gap in bilingual support and the need for stronger Chinese-language modeling in this domain.
Figure 6: Social media example with subtle bilingual inconsistency. The Chinese translation omits "student", changing the meaning and potentially misleading Chinese readers.
Failure Cases. We observe that BiMi occasionally fails when manipulations are extremely subtle or when OCR fails to extract accurate subtitles. In such cases, either the localization is off-target, or the explanation is overly general. More failure cases and analysis are provided in the Appendix.
6 Conclusion and Limitation
Conclusion. We study the task of detecting multimodal inconsistencies in news images with bilingual (Chinese-English) subtitles. To support this, we introduce BiMiBench, a large-scale benchmark with 104K samples featuring realistic visual edits and cross-lingual mismatches. We propose BiMi, a bilingual multimodal framework combining a three-stage post-training strategy and retrieval-augmented reasoning. Experiments show that BiMi achieves state-of-the-art performance in classification, localization, and explanation, highlighting its effectiveness in real-world scenarios.
Limitation and future work. While BiMi achieves strong results on bilingual multimodal detection, several limitations remain. First, the current setup focuses on Chinese-English subtitles; extending to more languages would enhance generalization. Second, localization is limited to bounding boxes, which may miss fine-grained manipulations—future work could explore pixel-level segmentation. Third, BiMiBench samples use fixed subtitle styles and editing templates; expanding visual diversity and subtitle formats would better reflect real-world cases. Finally, while BiMi offers solid performance, its reasoning and explanation ability can be further improved, especially in ambiguous scenarios.
Appendix
Section A provides details on the dataset, including: dataset statistics, misinformation generation strategy, example visualization, and manual verification.
Section B outlines implementation and experimental details, including: model architecture and training settings, hyperparameters.
Section C discusses the broader impact and ethical implications of our work.
Section D presents qualitative examples.
Section E analyzes the failure case.
Appendix A Dataset Details
Dataset Statistics. BiMiBench comprises 104,000 samples spanning six fine-grained misinformation categories. Figure 7 summarizes key dataset properties. The left panel shows the distribution of samples across categories, ensuring diverse coverage of manipulation types.
Figure 7: Dataset statistics of BiMiBench. Left: distribution across six misinformation categories. Center: English subtitle length distribution. Right: normalized area of manipulated regions (bounding boxes height).
Misinformation Generation Strategy. To simulate realistic multimodal misinformation, we generate manipulated subtitles by introducing subtle yet misleading semantic alterations. These edits preserve grammatical fluency and contextual plausibility while modifying key factual elements such as subjects, causal relations, locations, or numerical values. We further introduce bilingual inconsistencies by translating the manipulated English subtitles into Chinese, allowing semantic shifts to emerge naturally during translation. This strategy enables diverse and challenging misinformation scenarios that reflect real-world cross-modal and cross-lingual manipulations. We design task-specific prompts to guide misinformation generation, ensuring fluency and semantic plausibility while introducing factual inconsistencies; an example prompt used in this process is shown in Table 8.
<system> Based on the image, create a fictional, sensational news headline or summary in one sentence. The story must be clearly fake but sound plausible, and must be no more than 30 words. The news should sound outrageous, scandalous, or unbelievable, yet still written in a serious, news-like tone. </system>
Table 8: Prompt templates used in BiMi for multimodal misinformation detection
Example Visualization. Figure 8 illustrates representative BiMiBench samples, covering various manipulation types, language inconsistencies, and corresponding ground truth labels.
Figure 8: Sample BiMiBench examples with bilingual subtitles and visual or text manipulations.
Manual Verification. To ensure data quality and annotation reliability, all samples in BiMiBench were manually verified by a team of five trained annotators, each with backgrounds in journalism, linguistics, or media studies, and prior experience with fact-checking or misinformation analysis. Annotators received targeted instruction on identifying realistic image manipulations, assessing semantic consistency across English-Chinese subtitle pairs, and applying fine-grained manipulation categories. During the review process, they checked the accuracy of tampered regions, validated bounding box alignment, and evaluated bilingual subtitles for cross-lingual fidelity and plausibility. Samples with OCR errors, ambiguous edits, mistranslations, or low visual quality were discarded. This multi-layered, expert-driven verification process ensures that BiMiBench maintains high annotation quality, semantic precision, and supports reliable benchmarking for multilingual multimodal misinformation detection.
Appendix B Implementation Details
Model Architecture. Our framework is built upon the publicly released Gemma 3 multimodal language model. The vision encoder is initialized from a pretrained ViT-L model, and remains frozen throughout all training stages. LoRA parameters are fine-tuned. The model accepts the inputs of the image with English and Chinese subtitle, which serialized into a unified prompt format. The image is encoded through a unified visual-language backbone.
Tokenization and Input Format. Both English and Chinese subtitles are tokenized using the SentencePiece tokenizer aligned with the Gemma 3 backbone. The input is formatted into structured prompts that specify the prediction task, marking different modalities. Input length is capped at 512 tokens.
Hyperparameter Summary. We fine-tune BiMi using the Gemma 3 model (4B) with a batch size of 8, a learning rate of $1\times 10^{-5}$, and a maximum input length of 512 tokens. The model is optimized with AdamW, using $\beta_{1}=0.9$, $\beta_{2}=0.98$, and a weight decay of 0.01. To stabilize training, we use gradient accumulation (4 steps) and mixed precision (FP16). The GRPO reward coefficient $\lambda$s are all set to 1. These hyperparameters are kept consistent across all three training stages unless otherwise specified.
Prompt Template. We design a structured instruction format to guide the model in reasoning across modalities and generating faithful explanations. The prompt template used during instruction tuning and inference is shown in Table 9. This format ensures consistent output structure, supporting multimodal reasoning (via <think>), interpretable classification, and region-level manipulation localization.
<system> You are an expert in misinformation detection area. A conversation between User and Assistant. The user asks a question, and the Assistant solves it. The assistant first thinks about the reasoning process in the mind and then provides the user with the answer. The reasoning process and answer are enclosed within <think> </think> and <answer> </answer> tags, respectively, i.e., <think> reasoning process here </think><answer> answer here </answer>. Assign one of the following six categories based on their mutual alignment: All Consistent: The image aligns well with both Chinese and English captions. No signs of manipulation or misalignment. Image Manipulated: The image is manipulated. Both captions truthfully describe the manipulated image. Both Misaligned: Both Chinese and English captions are manipulated. Neither caption correctly describes the image. Chinese Misaligned: Only the Chinese caption is manipulated. The English caption aligns correctly with the image. English Misaligned: Only the English caption is manipulated. The Chinese caption aligns correctly with the image. All Inconsistent: The image, Chinese caption, and English caption are all manipulated and mutually inconsistent. </system> <user> Please analyze the given image containing both Chinese and English subtitles and complete the following three tasks: (1) Classification Task: classify the alignment between the image and the subtitles into one of the following six categories: "all consistent", "image manipulated", "both subtitles misaligned with image", "only English aligned", "only Chinese aligned", "all inconsistent". (2) Manipulation Detection: if the image has been manipulated, return one or more bounding boxes for the manipulated regions in the format: {"bbox":[x_min, y_min, x_max, y_max]}. If no manipulation is found, return an empty list. (3) Decision Explanation: briefly explain your thinking before the classification and any detected regions. Return your output using the following format, wrapped in tags: <think> Your explanation here. </think> <answer> "classification": "result", "region": [{"bbox": [x_min, y_min, x_max, y_max]}] </answer> </user>
Table 9: Prompt templates used in BiMi for multimodal misinformation detection
Appendix C Broader Impact and Ethics
BiMi is designed to enhance the interpretability and reliability of multimodal misinformation detection, particularly in the context of bilingual news media. By providing both classification outputs and natural language explanations, BiMi promotes transparency in AI decision-making and supports fact-checkers, journalists, and platform moderators in identifying subtle manipulations across images and subtitles. To ensure ethical standards, all data samples in BiMiBench have been manually reviewed, linguistic clarity, and annotation consistency. Low-quality, ambiguous content has been excluded. The dataset construction avoids personally identifiable information and sensitive content.
BiMi is detection-focused and does not produce or promote fake content, reducing the risk of misuse. Nonetheless, we recognize that any publicly released model or dataset could potentially be repurposed. To mitigate this, we include usage guidelines and encourage responsible deployment, especially in downstream applications that may involve high-stakes decisions. We believe this work contributes positively by advancing research on robust, interpretable AI tools for combating harmful information while maintaining strong ethical safeguards throughout the development and release process.
Appendix D Qualitative Examples
Figure 9 10 11 12 13 14 15 16 show the examples.
Figure 9: Case of All Consistency.
Figure 10: Case of Chinese Misaligned.
Figure 11: Case of Image Manipulation.
Figure 12: Case of English Misaligned.
Figure 13: Case of Both Misaligned.
Figure 14: Case of English Misaligned.
Figure 15: Case of All Inconsistent.
Figure 16: Case of Both Misaligned.
Appendix E Failure Cases
Despite the overall effectiveness of our model, we observed several notable failure cases that highlight its current limitations. One recurring issue involves semantic misinterpretation, particularly when the input contains rare or ambiguous phrases. For example, in a prompt requiring multi-step reasoning or contextual understanding (e.g., “Describe the implication of climate change on coastal microeconomies”), the model occasionally generates generic or overly broad responses that fail to capture the specific nuances of the query. Another type of failure arises in edge cases involving domain-specific knowledge.
Figure 17 shows a failure case where the model misclassified an image manipulation example as “All Inconsistent.” The English caption aligns with the image, but the Chinese one introduces a false espionage narrative. Key issues include: misclassification; incorrect caption judgment—only the Chinese caption was wrong; missing bounding box. This highlights weaknesses in manipulation detection, multilingual judgment, and grounding.
Figure 17: Failure case